OpenStudy (anonymous):

Will Medal And Fan Add. −4 + 4 =

OpenStudy (anonymous):

-4+4 can be written as (-4)+4 which in turn is same as 4+(-4)=4-4=0

OpenStudy (anonymous):

\(\large -4+4~=~4-4\)

OpenStudy (anonymous):

It is added. it says to add

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

@math1234

OpenStudy (anonymous):

anyone

OpenStudy (chrisdbest):

The answer is 0

OpenStudy (anonymous):

its adding

TheSmartOne (thesmartone):

Do what @LegendarySadist showed you. He just made the addition into a subtraction. It results in the same answer.

OpenStudy (chrisdbest):

Look: if you do -4 + 4 = 0! Just as 4 + -4 + 0!!!

OpenStudy (anonymous):

No. i put the answer in and it says no

OpenStudy (anonymous):

I can just look up the answer